A three-person maintenance crew could clean a certain building in 4h,whereas a four-person crew could do the job in 3 h. if one worker of the four person crew was an hour late , how long did the job take ?
HELP PLEASE !!!!!

Answer by richwmiller(17219)

3*4=12
4*3=12
One person does 1/12 an hour
In the first hour 3 workers do 3/12=1/4 of the job instead of 4/12 =1/3
Then 4 workers have to do 12-3=9 worker hours to finish
9/4=2 1/4 hours
So they take another 2.25 hours or a total of 3.25 hours